### v1

# Security modelling for cyber-physical systems: A systematic literature review

Abstract
Today, cyber-physical systems (CPS) find themselves at the intersection of digital technology and engineering domains, rendering them high-value targets of sophisticated and well-funded cybersecurity threat actors. Prominent cybersecurity attacks on CPS have brought attention to the vulnerability of these systems, and the soft underbelly of critical infrastructure reliant on CPS. Security modelling for CPS is an important mechanism to systematically identify and assess vulnerabilities, threats, and risks throughout system lifecycles, and to ultimately ensure system resilience, safety, and reliability.
This literature review delves into state-of-the-art research in CPS security modelling, encompassing both threat and attack modelling. While these terms are sometimes used interchangeably, they are different concepts. This article elaborates on the differences between threat and attack modelling, examining their implications for CPS security. A systematic search across leading scientific databases yielded 428 articles, from which 15 were selected and categorised into three clusters: those focused on threat modelling methods, attack modelling methods, and literature reviews. Specifically, we sought to examine what security modeling methods exist today, and how they address real-world cybersecurity threats and CPS-specific attacker capabilities throughout the lifecycle of CPS, which typically span longer durations compared to traditional IT systems.
This article also highlights several limitations in existing research, wherein security models adopt simplistic approaches that do not adequately consider the dynamic, multi-layer, multi-path, and multi-agent characteristics of real-world cyber-physical attacks. Against this backdrop, we propose a unified security modelling framework that aligns threat modelling, attack modelling, and security monitoring across the lifecycle of CPS. Last but not least, this article proposes future research directions in this evolving domain.
keywords:
cyber-physical systems, security modelling, threat modelling, attack modelling, systematic literature review, advanced persistent threats, self-healing systems, safety, reliability, resilience
1 Introduction
Cyber-physical systems (CPS) are complex systems that integrate physical processes with computational algorithms and network connectivity (Griffor et al. (2017)). These systems are designed to monitor, control, and optimize physical processes in real time, and are typically found in industrial automation, smart infrastructure, healthcare, and transportation domains. Today, CPS find themselves at the intersection of digital technology and engineering, and consequently become a high-value target of cybersecurity threat actors. Prominent cybersecurity attacks on Cyber-Physical Systems (CPS) (Kumar et al. (2022)) have brought attention to the vulnerability of these systems, and the soft underbelly of critical infrastructure reliant on CPS. These vulnerabilities could potentially result in significant consequences, affecting lives, economies, and national security.
Although CPS cybersecurity is often discussed in the context of critical infrastructure, CPS are increasingly interconnected in smart city infrastructure and the cloud. This presents cybersecurity risks not only from physical and cyber components, but also from less conventional but critical attack paths such as electric chargers. For example, Köhler et al. (2023) showed that the Combined Charging System, one of the most widely used DC rapid charging technologies for electric vehicles (EVs), could be exploited wirelessly to interrupt charging sessions for individual vehicles or entire fleets at the same time.
Today, CPS owners and operators face a daunting challenge in keeping pace with the increasing volatility, uncertainty, complexity, and ambiguity (VUCA) of cybersecurity threats. In IT systems, the consequences or impact of IT security incidents are more predictable and can usually be identified and mitigated before actual incidents occur. For example, companies can prepare holding statements to media, and purchase cybersecurity insurance to safeguard against reputational and financial loss before incidents occur. In contrast, CPS cybersecurity attacks are dynamic, less predictable, consequences may change quickly depending on the operating environment, and incident triage and investigations can be especially challenging because of the intricate and interconnected design of CPS.
Existing cybersecurity frameworks like the NIST Cybersecurity Framework (National Institute of Standards and Technology (2023)) which could help organisations manage and improve their cybersecurity risk management, and international standards like IEC 62443-4-1:2018 (International Electrotechnical Commission ) which specifies process requirements for the secure development of products used in industrial automation and control systems, offer valuable guidance for securing CPS. However, the dynamic nature and design constraints of CPS require a tailored approach, combining continual assessment and adaptation of physical measures, cybersecurity practices, as well as risk mitigation controls.
Two key elements of this approach are threat and attack modelling. Although both these terms are often used interchangeably, they are not the same. Threat modelling aims to anticipate cybersecurity threats—usually in the early stages of a system lifecycle, so that relevant mitigation controls can be incorporated into the system design (Uzunov and Fernández (2014); Dhillon (2011); Xiong and Lagerström (2019)). In contrast, attack modelling delves into specific attacker tactics, techniques, and procedures—which may not be limited to cyber but include physical actions as well—to develop targeted defences and mitigation strategies when the system is in operation (Al-Mohannadi et al. (2016); Saini et al. (2008); Schneier (1999); Zenitani (2023)). On this note, threat and attack modelling need to be tailored for CPS and can be an important tool for critical infrastructure owners and operators when navigating cybersecurity threats faced in CPS (Yang and Zhang (2023); Paudel et al. (2017)).
Contributions. In this review, we identified a number of state-of-the-art threat modelling and attack modelling studies relevant to cyber-physical systems (CPS), with the following findings:
  • Threat modelling is commonly conducted in early stages of system development. However, this means that evolving attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles will render threat models outdated and irrelevant.
  • Security models in the literature generally focus on IT systems, and these models are challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • Most papers do not differentiate between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks, effects and consequences are considered in security modelling.
  • There is ambiguity in the literature regarding the definitions and relationship between threat modelling and attack modelling. Correspondingly, a unified security modelling framework that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article.
The rest of the article is structured as follows: Section 2 provides an overview of the review’s background while Section 3 describes related work. Our methodology and results are outlined in Sections 4 and 5. Based on the results, we discuss our findings and research directions for future work in Section 6, and conclude in Section 7.
2 Background
This section provides some background into how CPS differ from conventional IT systems, and the characteristics of adversaries who target CPS in cyber attacks. Together, this explains why a tailored security modelling approach is necessary for CPS.
2.1 Cyber-Physical Systems
Different from traditional IT systems, a CPS is a System of Systems (SoS) that integrates computational elements, communication networks, and physical processes to form a unified system. CPS converges Information Technology (IT) and Operational Technology (OT) and has different priorities for cybersecurity. OT focuses on safety and reliability, while IT emphasises the confidentiality, integrity, and availability of information. Additionally, CPS networks must accommodate diverse communication modes, ranging from standalone to highly networked systems, potentially utilising legacy protocols such as serial communications, TCP/IP, or object exchange protocols. The heterogeneity and, in certain instances, expansive geographic scopes of CPS networks contribute to the complexity of modelling security for CPS.
Despite their longer system lifecycles, CPS are not upgraded or patched frequently as system or software changes may affect their reliability or even safety (Kavallieratos et al. (2020); Kriaa et al. (2015); Sabaliauskaite and Mathur (2014); Suo et al. (2018)). This results in outdated software, exposing CPS to an ever-increasing number of cybersecurity vulnerabilities over time. At the same time, some CPS employ “security-by-obscurity”, using proprietary software and technology that may contain undisclosed security weaknesses that attackers could exploit. Moreover, the lack of “security-by-design” in CPS lowers the resilience to cybersecurity attacks even further, as CPS may lack security logs and fail to leverage log data for proactive cybersecurity monitoring and defence. Finally, there is a misconception that merely physically isolating Cyber-Physical Systems (CPS) from external networks, known as an “air gap” network security measure, is adequate to shield CPS from all cybersecurity threats. However, this notion is flawed, particularly in scenarios where multi-agent, multi-path CPS attacks could employ social engineering techniques, physical access, insider threats, and exploit supply-chain vulnerabilities, enabling attackers to breach the “air gap” and compromise CPS, even when they are physically segregated from external networks.
2.2 Adversarial model
MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) (Strom et al. (2018)) serves as a widely adopted knowledge base for characterising the actions and behaviours of cyber adversaries. Specifically tailored for Industrial Control Systems (ICS)—a category of Cyber-Physical Systems (CPS)—MITRE ATT&CK provides a matrix known as ”ATT&CK for ICS” (Alexander et al. (2020)). This matrix focuses on the unique Tactics, Techniques, and Procedures (TTPs) employed by adversaries targeting industrial control environments. These TTPs play a crucial role in informing and guiding cybersecurity defence for CPS across the various stages of a cyber-attack lifecycle, as outlined in the Lockheed Martin cyber kill chain model (Hutchins et al. (2011)).
An analysis of past cyber attacks on CPS reveals a taxonomy (Zhu et al. (2011)) highlighting differences from those observed in IT systems, necessitating a tailored security modelling approach for CPS. CPS attackers employ dynamic, multi-agent, and multi-path attack TTPs to defeat physical access controls and “airgap” measures (isolating CPS from external networks). Additionally, adversaries evolve their TTPs over time, such as using “low and slow” TTPs to evade detection and prolong access to compromised CPS.
2.3 CPS cyber-intrusion example
We present an example of a 6-month long CPS intrusion by a sophisticated attacker group comprising two adversaries A and B, adapted from one of the author’s professional experience. In this example, Adversary A compromises the organisation’s IT network to gain remote access to a CPS engineer’s computer and employs social engineering techniques to deliver custom malware to an isolated (“air gapped”) CPS network. Adversary B’s focus on the other hand is to develop custom malware to firstly profile the CPS, then to weaponise custom malware to manipulate connected Remote Terminal Units (RTUs) causing physical damage e.g. via actuators or valves. The diamond model of intrusion analysis (Caltagirone et al. (2013)), as well as a combination of the diamond model and the cyber kill-chain (Ertaul and Mousa (2018)) are used to describe both the flow (Figure 1) and timeline (Table 1) of this CPS cyber-attack.
Figure 1: CPS cyber intrusion flow
Table 1: CPS cyber intrusion timeline
Phase
January
February
March
April
May
June
Reconnaissance
Scan for email servers, find account credentials
Find users or systems that have access to CPS
Weaponisation
Customise malware for CPS information gathering
Customise malware for RTU reset
Delivery
Deploy malware to CPS through engineer’s USB thumb-drive
Deploy malware to CPS through engineer’s USB thumb-drive
Exploitation
Access VPN
Compromise CPS engineer’s computer
Suppress alarm reporting to main server
Scheduled task manipulates connected RTUs
Installation
Install Cobalt Strike beacon agent
Install Cobalt Strike beacon agent
Malware runs and collects network information
Malware runs on CPS and hides process, awaits scheduled task
C2
Establish C2 callback from IT network
Establish C2 callback
Maintain access
CPS information saved in USB thumb-drive
Maintain access
Maintain access
Actions on Objectives
Services disrupted
2.4 Security modelling framework for CPS
In the CPS cyber intrusion example, we demonstrated the key characteristics and inter-relationships between threat and attack modelling methods, namely the diamond model and cyber kill-chain respectively. Against this backdrop, we propose a unified security modelling framework for CPS (Figure 2) that integrates these modelling methods while drawing from continuous security monitoring across all stages of the CPS lifecycle for informed cybersecurity decision-making.
This security modelling framework for CPS does not replace threat modelling and attack modelling methods. Rather, the framework makes it easier and more practical for CPS cybersecurity practitioners to adopt and use these modelling methods to address evolving cybersecurity threats and attacker capabilities throughout the CPS lifecycle.
Figure 2: Security modelling framework for CPS
3 Related Work
There are few existing systematic literature reviews (SLR) on threat modelling or attack modelling. Humayed et al. (2017)’s literature survey covered the security and privacy of cyber-physical systems, with a special focus on ICS, smart grids, medical devices, and smart cars. They presented a taxonomy of CPS threats, vulnerabilities, known attacks and existing controls, and importantly, captured how an attack of the physical domain of a CPS can result in unexpected consequences in the cyber domain and vice versa.
The SLR by Xiong and Lagerström (2019) was another early review on cybersecurity threat modelling, where one CPS threat modeling method was cited: Burmester et al. (2012)’s framework. Based on the Byzantine paradigm, it enables the modeling of CPS security and facilitates formal analysis and security proof through cryptographic methodologies. Tatam et al. (2021) reviewed threat modelling approaches for APT-style attacks, but the review did not focus on CPS security. The SLR by Khalil et al. (2023) on the other hand, focused on threat modelling of industrial control systems (ICS). The authors described the various ICS threat models adopted in literature, and noted it was timely to consider a framework that covers all aspects of the various threat modelling methods.
Earlier, the evaluation of threat modelling methods by Shevchenko et al. (2018) made a similar conclusion that there is no one method that can cover the full spectrum of CPS threats, and a framework that employed a combination of methods and techniques should be used. Supporting this argument, the hybrid threat model for smart systems proposed by Valenza et al. (2023), and the work by Li et al. (2021) on security-usability threat modelling for industrial control systems underscored the need to consider not only cyber but also human factors in ICS threat modelling due to the system-of-systems (SoS) nature of CPS.
These CPS threat modelling methods and techniques were well-covered in the literature, with case studies in specific industries and domains, e.g. the study of smart grid CPS cybersecurity by Nafees et al. (2023), the study of cyber-physical energy systems security (CPES) by Zografopoulos et al. (2021), the application of the STRIDE threat model in modern vehicle vulnerability assessments by Abuabed et al. (2023), the work on CPS threat modelling for power transformers by Ahn et al. (2021), the work on medical CPS threat modelling by Almohri et al. (2017), the work by Al Asif and Khondoker (2020) on telesurgery system threat modelling, the work by Jbair et al. (2022) on threat modelling for industrial cyber-physical systems (ICPS) in the manufacturing industry, and the work on threat modelling in smart firefighting systems by Zahid et al. (2023).
4 Methodology
We followed the guidelines for systematic literature reviews proposed by Kitchenham and Charters (2007). The process of our literature review is shown in Figure 3.
Figure 3: Literature review process
4.1 Research questions
This review aims to answer the following Research Questions (RQs):
  • RQ1: What cyber-physical system (CPS) threat or attack models have been adopted in existing literature?
    The rationale for this research question is to understand what the “state-of-the-art” research in this area is.
  • RQ1.1: How do the proposed threat and attack models align with cybersecurity in various phases of the cyber-physical system lifecycle?
    The rationale for this research question is to understand how existing CPS threat and attack models address cybersecurity threats not only in the design phase, but when CPS are in operational phases. This is particularly crucial as CPS may operate for many years, and likely render threat models considered in the system design irrelevant.
  • RQ1.2: In what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?
    The rationale for this research question is to understand how existing CPS threat and attack models are augmented and improved with real-world feedback from cybersecurity intrusions, near-misses, or external threat intelligence information. This is an important bridge between theoretical models and practice, that can lead to tangible security improvements in CPS.
  • RQ2: What research gaps have been identified in the methods used for modelling threats and attacks in cyber-physical systems (CPS)?
    The rationale for this research question is to help identify research gaps in current CPS security modelling methodologies and to direct future research efforts.
4.2 Search protocol
This sub-section aims to describe the search protocol that we used in this review.
We used US English for all search terms. For example, we used “modeling” instead of “modelling”, in UK English. We validated that Scopus did not distinguish between the spellings and returned the same search results (Figure 4).
Figure 4: Comparison of search results between US and UK English keywords
We selected IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus to perform our search for articles. These scientific databases were chosen as they are recognised as reputable and accurate repositories for academic literature, and they have a good collection of results relating to cybersecurity articles and publications.
For the search keywords, we were primarily interested to search: cybersecurity modelling, security modelling, threat modelling and attack modelling, in line with the subject of our literature review. As such, we specified our first search keywords as: “security model”, “security modeling”, “cybersecurity model”, “cybersecurity modeling”, “threat model”, “threat modeling”, “attack model”, and “attack modeling”. We included both words, “model” and “modeling” to increase the coverage of search results.
As we wanted to narrow our search results to articles on security, threat, or attack modelling, we limited the initial search to article titles. There was a need to make minor changes to the search terms, depending on the Boolean vocabulary used on the database.
  • On IEEE Computer Society Digital Library, we selected Advanced Search, selected the “in Document Title” option, and entered the search keywords: (model modeling) AND (cybersecurity security threat attack).
  • On ScienceDirect, we selected Advanced Search and entered (“cybersecurity model” OR “security model” OR “threat model” OR “attack model”) in the Title field. The Boolean “OR” had to be specified, and there was no need to specify the “modeling” keyword as it is implied by “model” in the search term.
  • On ACM Digital Library, we selected “Edit Search”, selected “Title” under “Search Within”, then entered: (“security model” “security modeling” “cybersecurity model” “cybersecurity modeling” “threat model” “threat modeling” “attack model” “attack modeling”).
  • On Scopus, we selected “Article title” under the “Search within” field, then entered “security model” OR “security modeling” OR “threat model” OR “threat modeling” OR “attack model” OR “attack modeling” OR “cybersecurity model” OR “cybersecurity modeling” in the “Search documents” field.
We then refined the search results further by specifying additional keywords: (“cyber physical” CPS SCADA ICS APT), to search within article titles, abstracts and contents. We included SCADA (Supervisory Control and Data Acquisition), ICS (Industrial Control System) and APT (Advanced Persistent Threat) in the search, so that search results would be relevant to the adversarial tactics, techniques and procedures associated with CPS cyber attacks.
  • On IEEE Computer Society Digital Library, we used the Advanced Search menu with the previous keywords, selected the “AND” and “in Anything” options, and entered the keywords: (“cyber physical” CPS SCADA ICS APT).
  • On ScienceDirect, we selected Advanced Search and entered (“cyber physical” OR CPS OR SCADA OR ICS OR APT) in the “Find articles with these terms” field.
  • On ACM Digital Library, we selected “Edit Search”, added a new search field, selected “Anywhere” under “Search Within”, then entered: (“cyber physical” OR CPS OR SCADA OR ICS OR APT).
  • On Scopus, we selected “Add search field”, selected “Article title, Abstract, Keywords” under the “Search within” field, and entered: “cyber physical” OR CPS OR SCADA OR ICS OR APT.
In the final phase of the search protocol, we removed repeated articles from the combined search results, based on similar article titles. We also removed articles that were not related to the research questions, after reviewing article titles and abstracts.
4.3 Categorisation and classification
We categorised the selected articles into three clusters according to the following criteria:
  • Cluster 1 (C1): Article focuses on threat modelling method(s)
  • Cluster 2 (C2): Article focuses on attack modelling method(s)
  • Cluster 3 (C3): Article is a systematic literature review (SLR)
Additionally, we classified the articles according to the following criteria:
  1. 1.
    Year of publication
  2. 2.
    Document type, i.e. conference paper, article, review, book chapter, etc.
  3. 3.
    Number of citations (based on Google Scholar Citations)
  4. 4.
    Research question(s) article is relevant to
4.4 Quality assessment
In this phase, we reviewed the articles to ensure they were relevant to CPS security, threat or attack modelling, and we removed articles that do not describe these models’ approaches, CPS applications, or CPS case-studies. We also reviewed bibliographies in the articles and included additional articles that may be relevant to the review.
4.5 Data collection
We studied the articles closely to understand the background, approaches, and results from the various research. Importantly, we sought to identify gaps in the research in terms of security modelling in the context of CPS cyber attacks. We documented our notes from the review in a shared repository (Huang (2024)) for future reference.
4.6 Data analysis
In the final phase of the SLR methodology, we synthesised our findings and documented our analysis as well future research directions.
5 Results
The search was conducted on the following digital libraries in January 2024: IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus. Following the search, we had 438 articles in total. We performed a rigorous quality assessment, where we first excluded duplicate articles, then we identified and excluded irrelevant articles based on their abstracts. We then performed full-reading of about 27 articles that remained and excluded articles that were not relevant to any of our research questions. Finally, we had 15 articles subject to further analysis (Figure 5).
Figure 5: Search results (January 2024)
The articles were categorised into three clusters (Table 2) based on full reading of the articles: threat modelling (C1), attack modelling (C2), and SLRs (C3). There were no articles on “security modelling” that were relevant to CPS.
Table 2: Categorisation of articles
Reference
Title
Threat Modelling (C1)
Attack Modelling (C2)
SLRs (C3)
Ayrour et al. (2018)
Modelling cyber attacks: a survey study
$\bullet$
Fernandez (2016)
Threat modeling in cyber-physical systems
$\bullet$
Khalil et al. (2022)
Threat Modeling of Cyber-Physical Systems-A Case Study of a Microgrid System
$\bullet$
Khalil et al. (2023)
Threat modeling of industrial control systems: A systematic literature review
$\bullet$
Khan et al. (2017)
STRIDE-based threat modeling for cyber-physical systems
$\bullet$
Kim et al. (2022)
STRIDE‐based threat modeling and DREAD evaluation for the distributed control system in the oil refinery
$\bullet$
Kumar, R., et al. (2022)
APT attacks on industrial control systems: A tale of three incidents
$\bullet$
Martins et al. (2015)
Towards a systematic threat modeling approach for cyber-physical systems
$\bullet$
Mekdad et al. (2021)
A threat model method for ICS malware: the TRISIS case
$\bullet$
Neubert and Vielhauer (2020)
Kill chain attack modelling for hidden channel attack scenarios in industrial control systems
$\bullet$
Paudel et al. (2017)
Attack models for advanced persistent threats in smart grid wide area monitoring
$\bullet$
Tatam et al. (2021)
A review of threat modelling approaches for APT-style attacks
$\bullet$
Xiong and Lagerström (2019)
Threat modeling – A systematic literature review
$\bullet$
Yang and Zhang (2023)
From Tactics to Techniques: A Systematic Attack Modeling for Advanced Persistent Threats in Industrial Control Systems
$\bullet$
Zalewski et al. (2013)
Threat modeling for security assessment in cyberphysical systems
$\bullet$
There were 10 articles in C1 and five in C2, showing that existing literature delved into threat modelling methods more than attack modelling methods. Next, we then classified the articles based on the criteria described in Section 4.3.
The first selected article was published in 2013 and although no article was published in 2014, the number of relevant articles has gradually risen over the years (Figure 6), signaling a growing level of research interest in CPS security modelling. Furthermore, out of the 15 selected articles, seven were from conference proceedings and eight were from journals. This indicated that almost an equal number of articles in CPS cybersecurity modelling were published in journals and conferences.
2014201520162017201820192020202120222023111122223333
Figure 6: Number of articles per year
Next, we used citation metrics on Google Scholar to identify articles that had more than 100 citations. As of January 2024, only two of the 15 selected articles had more than 100 citations (Table 3). The top cited article was on STRIDE-based threat modelling for cyber-physical systems by Khan et al. (2017), followed by the SLR on threat modelling by Xiong and Lagerström (2019). This observation, together with that from the number of articles published yearly, suggested that this field of research is still relatively new.
Table 3: Top-cited articles
Reference Citations
Khan et al. (2017) 239
Xiong and Lagerström (2019) 231
Finally, we performed full reading of the 15 articles and analysed which research questions they were relevant to (Table 4).
Table 4: Article relevance to research questions
Reference Cluster RQ1 RQ1.1 RQ1.2 RQ2
Fernandez (2016) C1 $\bullet$
Khalil et al. (2022) C1 $\bullet$ $\bullet$
Khan et al. (2017) C1 $\bullet$
Kim et al. (2022) C1 $\bullet$ $\bullet$
Martins et al. (2015) C1 $\bullet$
Mekdad et al. (2021) C1 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Zalewski et al. (2013) C1 $\bullet$ $\bullet$
Kumar et al. (2022) C2 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Neubert & Vielhauer (2020) C2 $\bullet$ $\bullet$ $\bullet$
Paudel et al. (2017) C2 $\bullet$ $\bullet$ $\bullet$
Yang and Zhang (2023) C2 $\bullet$ $\bullet$ $\bullet$
Ayrour et al. (2018) C3 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Khalil et al. (2023) C3 $\bullet$ $\bullet$
Tatam et al. (2021) C3 $\bullet$ $\bullet$
Xiong and Lagerström (2019) C3 $\bullet$
Three of the articles (Ayrour et al. (2018); Kumar et al. (2022); Mekdad et al. (2021)) addressed all the research questions. The remaining articles addressed some, but not all the research questions. 12 of the 15 articles were relevant to RQ1, and only two of the seven C1 articles were relevant to RQ1.1 (28.57%), one was relevant to RQ1.2 (14.29%), and three (42.86%) were relevant to RQ2. On the other hand, three of the four C2 articles were relevant to RQ1.1 (75%), four were relevant to RQ1.2 (100%) and two were relevant to RQ2 (50%). Lastly, three of the four C3 articles were relevant to RQ1.1 (75%), one was relevant to RQ1.2 (25%), and four were relevant to RQ2.
A comparison of articles across the clusters (Figure 7) showed that existing literature that focused on threat modelling (C1) addressed RQ1 more than the other research questions. Those that were focused on attack modelling (C2) appeared to address RQ2 less. Lastly, the C3 articles did not address RQ1 as they were literature reviews, and these articles appeared to address RQ1.2 less.
C1C2C30050505050100100100100ClusterRelevant articles (%)RQ1RQ1.1RQ1.2RQ2
Figure 7: Relevant articles across clusters
6 Discussion
This section aims to describe the findings and limitations of this review, as well as future research directions.
6.1 Findings
The selected 15 articles were categorised into three clusters: threat modelling methods (C1), attack modelling methods (C2), and SLRs (C3). There were eventually seven C1 articles, five C2 articles, and three C3 articles.
6.1.1 RQ1
To answer the first research question, “what cyber-physical system (CPS) threat or attack models have been adopted in existing literature?”, we aimed to identify the threat and attack models that were adopted in the selected articles. This was not straightforward as authors appeared to have varying definitions of threat and attack modelling methods, and SLRs in particular, categorised both as threat modelling methods.
For example, Khalil et al. (2023) proposed “attack-centric threat modelling” in their ICS threat modelling taxonomy, and distinguished the term from attack modelling, stating that the latter focused solely on the attacker’s behaviour and not the defender’s perspective. In addition, Tatam et al. (2021) categorised models that are typically used to analyse cyber attacks, such as the diamond intrusion analysis model (Caltagirone et al., 2013), attack trees (Saini et al. (2008); Schneier (1999)), the kill chain (Hutchins et al. (2011)), and the MITRE ATT&CK framework (Strom et al. (2018)) as threat models.
As such, we used the clusters as a basis for identifying threat and attack modelling methods described in the articles (Table 5), since the categorisation was based on research objectives and approach. We found that this approach produced a reasonable result as the modelling methods appeared to be categorised correctly. One exception was the article by Mekdad et al. (2021) where the diamond intrusion analysis model may be more accurately categorised as an attack modelling method.
Table 5: Threat and attack modelling methods adopted in selected articles
Reference Cluster
Threat Modelling Method(s)
Attack Modelling Method(s)
Fernandez (2016) C1
Misuse patterns
Khalil et al. (2022) C1
STRIDE
Khan et al. (2017) C1
STRIDE
Kim et al. (2022) C1
STRIDE, DREAD
Martins et al. (2015) C1
Generic Modeling Environment (GME)
Mekdad et al. (2021) C1
Diamond Model of Intrusion Analysis
Zalewski et al. (2013) C1
STRIDE
Kumar et al. (2022) C2
Attack tree
Neubert and Vielhauer (2020) C2
Kill chain
Paudel et al. (2017) C2
Attack tree
Yang and Zhang (2023) C2
Abstract APT attack model
6.1.2 RQ1.1
On the second research question, “how do the proposed threat and attack models align with cybersecurity in various phases of the cyber-physical system lifecycle?”, we found that threat modelling was conducted in the early stages of the system development cycle (Khalil et al. (2022)), or more specifically, during system design and validation phases (Khan et al. (2017); Kim et al. (2022); Martins et al. (2015)). Interestingly, unlike threat modelling, none of the C2 articles suggested or mentioned which phase of the system lifecycle attack modelling should be performed in. A possible reason is that attack modelling is seen as a complement to threat modelling methods (Khalil et al. (2022); Kumar et al. (2022)).
6.1.3 RQ1.2
On the third research question, “in what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?”, we found that there were different views on incorporating attacker tactics, techniques, and procedures (TTPs) into threat modelling frameworks. For example, Khalil et al. (2022) viewed that attack categories in their proposed attack taxonomy should be abstract and should not include details of the attack techniques such as those in MITRE ATT&CK for Industrial Control Systems (Alexander et al. (2020)). In contrast, Fernandez (2016) proposed a threat model using misuse patterns which may include attacker TTPs to represent CPS threats.
Likewise, several articles made references to attacker TTPs, or similar attacker behavioural characteristics in their analysis. Mekdad et al. (2021) combined the Diamond Model of intrusion analysis (Caltagirone et al. (2013)) and the ICS kill chain (Assante and Lee (2015)) to generate the ICS threat model. Attack trees (Schneier (1999)) were used to model CPS attacks (Kumar et al. (2022); Paudel et al. (2017)), while Neubert and Vielhauer (2020) presented how the Lockheed Martin Cyber Kill Chain (Hutchins et al. (2011)) can be used to model hidden channel attack scenarios in ICS. Yang and Zhang (2023) proposed an abstract APT attack model that can recognise and correlate attacker TTPs to identify and mitigate complex attacks in ICS, highlighting that the need to consider the range of TTPs used by sophisticated adversaries to represent real threats and reflect the evolving threat landscape.
Importantly, we noted that while several articles considered the need for threat or attack models to account for changes over time, they did not include attacker TTPs as one of those changes. For example, Almohri et al. (2017) recognised that the threat model of a medical CPS is related to its corresponding trust model, and the trust model may include time-bound elements such as temporarily-trusted individuals or accounts. Additionally, Khalil et al. (2022) noted that systems may change over time and information initially excluded from threat analysis should be included in future threat models to account for system modifications.
Our findings raise the question of whether threat modelling continues to be relevant in CPS security design if they are conducted in the early stages of system development, given the usually long periods of development and implementation for complex CPS infrastructure. The cyber threats and available technologies to defend against those threats would have changed by the time the CPS become operational, often resulting in insecure systems. Relatedly, this presents a critical issue potentially affecting national security. CPS are usually used in critical national infrastructure and for that reason, are a prime target of sophisticated adversaries who have extended resources, capabilities and experience, may adopt “low and slow” tactics to infiltrate and maintain persistence in compromised CPS over many months and years. Outdated threat models would not be able to adjust and enable swift detection and eradication of these persistent threats.
6.1.4 RQ2
For our final research question, “what research gaps have been identified in the methods used for modelling threats and attacks in cyber-physical systems (CPS)?”, we made several observations from the literature.
Security models mostly focus on IT systems. Firstly, threat and attack models in the literature mostly focus on IT systems, even in the context of CPS security. For example, the STRIDE model developed by Microsoft was used in several articles to identify CPS threats, notwithstanding that the model is more commonly used in the context of software development and system security. Using STRIDE in CPS threat modelling (Abuabed et al. (2023); Khan et al. (2017); Kim et al. (2022)) has several limitations which were not obvious in the literature.
Specifically, the STRIDE approach relies on Data Flow Diagrams (DFDs) to identify system entities and trust boundaries, which has several shortcomings related to inadequate representation of security concepts, data elements, abstraction levels, and deployment information (Sion et al. (2020)). Moreover, it is sometimes not practical or even possible to construct an accurate DFD in CPS. There are various reasons for this: a CPS is a system-of-systems that includes both cyber and physical components; availability, accuracy and integrity of data flows between cyber and physical components cannot be assumed; and most importantly, the consequences of CPS compromise extend beyond those specified in the STRIDE model, e.g. physical changes or effects.
This limitation may be addressed by leveraging diverse, multi-disciplinary expertise from physical security and domain specialists to augment the DFD by capturing interdependencies and interactions between cyber and physical components, and not be limited to IT systems; and by continuously refining the DFD based on real-time data and feedback from system operations and cybersecurity events detected in the CPS environment over time.
Relatedly, the literature refers to CPS use-cases in the context of industrial control systems (ICS), smart grids, medical CPS (MCPS), supervisory control and data acquisition (SCADA) systems, or modern vehicles. We expect that CPS use-cases will increasingly include new scenarios involving technologies such as robotics, drones, EV charging infrastructure and smart buildings. Conventional threat modelling approaches, primarily tailored for IT systems, may encounter heightened challenges when addressing cybersecurity threats in these emerging domains.
Attack scenarios are not realistic. Secondly, most threat and attack modelling methods in the literature made assumptions based on simplistic cybersecurity attack scenarios, and did not consider the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS. The exceptions were: Fernandez (2016) who proposed a patterns-based threat model that considers multiple vulnerabilities exploited in different parts of a distributed system, and Zografopoulos et al. (2021) who analysed Cyber-Physical Energy Systems (CPES) security from a multi-layered attack perspective.
Differentiation between IT and CPS cybersecurity breaches. Thirdly, most of the literature did not differentiate cybersecurity breaches in CPS from those in IT systems in that they are not a binary (all or nothing) event. Instead of abrupt, catastrophic system failures, CPS cybersecurity breaches often result in gradual degradation of system services before the CPS fails eventually (Zalewski et al. (2013)). This is analogous to a system fault in an autonomous vehicle, the vehicle does not come to a stop immediately—which would injure the passengers—but instead gradually slows to a safe stop. The implication to CPS security is that from a cyber defence perspective, there is an opportunity to remediate intrusions through cyber or physical interventions (Ertaul and Mousa (2018)) or to perform dynamic reconfiguration and self-healing tasks (Zografopoulos et al. (2021)) before the CPS fails completely.
Consequence-driven considerations. CPS increasingly face cybersecurity threats that can endanger the safety, reliability and resilience of critical infrastructure. With the convergence of cybersecurity, safety, and reliability in CPS, it would be beneficial to adopt a consequence-driven and cyber-informed approach (Freeman et al. (2016)) towards threat and attack modelling, and align with system development processes in engineering domains such as the V-model that is commonly used in the automotive industry (Bolz et al. (2020)), as well as the failure mode and effect analysis (FMEA) and fault-tree analysis (FTA) techniques used in safety and reliability engineering domains (Kriaa et al. (2015)). Doing so will avoid making threat or attack modelling a one-time event that assumes static cybersecurity threats and vulnerabilities, resulting in inherent security weaknesses that are “baked” into proprietary CPS systems and software designs.
Distinguishing between CPS threat and attack modelling. Lastly, we found that while the literature used “threat modelling” and “attack modelling” interchangeably, both terms usually referred to threat modelling and not the latter. On the other hand, some authors promoted definitions of threat modelling from literature and distinguished threat from attack modelling (Khalil et al. (2023); Xiong and Lagerström (2019)), while others proposed security modelling frameworks that are based on threat models (Burmester et al. (2012)). Based on the literature, we analysed the various characteristics and inter-relationships between threat and attack modelling methods and these are aligned with our proposed unified security modelling framework for CPS as described in Section 2.4.
6.2 Future work
The limitations identified in the literature can be summarised as follows: existing security models predominantly focus on IT systems, neglecting the unique complexities of CPS scenarios and overlooking the multi-layered, multi-path, and multi-agent nature of real-world CPS attacks. Additionally, they often do not distinguish the consequences of cybersecurity breaches in CPS from those in IT systems and tend to treat cybersecurity threats as static and one-time events, contributing to inherent security weaknesses embedded within proprietary CPS systems and software designs.
Against this backdrop, our future work may be to research cybersecurity anomaly detection in CPS, to review the extent to which real-world cybersecurity intrusions and threat intelligence can inform and at the same time, enhance security modelling throughout the CPS system lifecycle. A more context-driven security model as this would better enable the identification of vulnerabilities, understanding of attack propagation mechanisms, and evaluation of system resilience.
We may also use Petri net based approaches (Chen et al. (2011); Dahl and Wolthusen (2006)) to better model CPS intrusions, integrating both cyber and physical actions. Petri nets are a graphical and mathematical tool for studying concurrent and distributed systems and have been widely applied in many different areas of computer science and other disciplines (Murata (1989)). In the context of CPS intrusion analysis, Petri nets are invaluable for modelling the interactions between cyber and physical components over time, thereby overcoming limitations of existing security modelling approaches when considering dynamic, multi-agent cyber-physical threats.
Finally, our future work may be to research self-healing techniques in CPS, to bolster existing cybersecurity defences and strengthen system resilience whether through cyber or physical actions, against sophisticated adversaries and threats.
6.3 Limitations of this review
Although we aimed to cover as much existing literature on security, threat and attack modelling in CPS as possible, we limited our search protocol to leading scientific databases as the quality of the articles may be ensured. Furthermore, we limited our search to journals and conference proceedings and consequently, we may miss out relevant articles from other sources such as book chapters, etc.
We also limited the initial scope of search to keywords in the article titles. This approach may have excluded articles relevant to security modelling, threat modelling, or attack modelling that did not explicitly mention these terms in their titles. However, to partially mitigate this exclusion, we reviewed the bibliographies of the selected papers and added any other relevant articles that may have been missed initially. Nevertheless, we ensured to select only articles that are pertinent to the research questions for eventual review. Based on the review of abstracts of the shortlisted articles from the search, we believe the coverage of the review to be sufficiently comprehensive.
Finally, our research questions were focused on how existing security modelling approaches addressed CPS, and corresponding adversarial tactics, techniques, and procedures. These factors may introduce some bias to the articles selected for review and may miss out articles that are relevant to CPS but did not mention adversarial models specifically. We addressed this by reviewing bibliographies in the selected articles and then including additional articles that may be relevant to the review.
7 Conclusion
In this review, we identified a number of state-of-the-art threat modelling and attack modelling studies relevant to cyber-physical systems (CPS), with the following findings.
  • Threat modelling is commonly conducted in early stages of system development. However, this means that evolving attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles will render threat models outdated and irrelevant.
  • Security models in the literature generally focus on IT systems, and these models are challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • Most papers do not differentiate between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks, effects and consequences are considered in security modelling.
  • There is ambiguity in the literature regarding the definitions and relationship between threat modelling and attack modelling. Correspondingly, a unified security modelling framework that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article.
While the number of published articles is comparatively small compared to more established research domains, there is visible growing interest in CPS cybersecurity research, as evidenced by the rising number of publications since 2013. Anticipating the broader use of CPS beyond traditional industrial scenarios, we envisage continued improvements in both the diversity and quality of articles in this domain, underscoring the growing significance of CPS cybersecurity.