### v2

# Security Modelling for Cyber-Physical Systems: A Systematic Literature Review

Abstract
Cyber-physical systems (CPS) are at the intersection of digital technology and engineering domains, rendering them high-value targets of sophisticated and well-funded cybersecurity threat actors. Prominent cybersecurity attacks on CPS have brought attention to the vulnerability of these systems, and the inherent weaknesses of critical infrastructure reliant on CPS. Security modelling for CPS is an important mechanism to systematically identify and assess vulnerabilities, threats, and risks throughout system lifecycles, and to ultimately ensure system resilience, safety, and reliability. This survey delves into state-of-the-art research in CPS security modelling, encompassing both threat and attack modelling. While these terms are sometimes used interchangeably, they are different concepts. This article elaborates on the differences between threat and attack modelling, examining their implications for CPS security. A systematic search yielded 428 articles, from which 15 were selected and categorised into three clusters: those focused on threat modelling methods, attack modelling methods, and literature reviews. Specifically, we sought to examine what security modelling methods exist today, and how they address real-world cybersecurity threats and CPS-specific attacker capabilities throughout the lifecycle of CPS, which typically span longer durations compared to traditional IT systems. This article also highlights several limitations in existing research, wherein security models adopt simplistic approaches that do not adequately consider the dynamic, multi-layer, multi-path, and multi-agent characteristics of real-world cyber-physical attacks.
keywords:
Cyber-physical systems, security modelling, threat modelling, attack modelling, systematic literature review, advanced persistent threats, self-healing systems, safety, reliability, resilience
1 Introduction
Cyber-physical systems (CPS) [1] are facing increasingly unconventional cybersecurity threats. Unlike traditional cybersecurity threats, modern cybersecurity threats are volatile, uncertain, complex, and ambiguous (VUCA). These threats are exacerbated by the wide interconnectivity of CPS in smart city infrastructure and the cloud. This interconnectivity presents a myriad of cybersecurity risks from both physical and cyber domains, including unconventional attack paths. In one example of an unconventional attack path, Kohler et al. [2] demonstrated that the Combined Charging System, a widely used DC rapid charging technology for electric vehicles (EVs), could be exploited wirelessly to interrupt charging sessions for individual vehicles or entire fleets at the same time.
The severity of CPS cyber-attacks stem from the roles that CPS perform in monitoring, controlling, and optimising critical operations that serve cities and communities. The critical role of CPS further underscores the need to ensure they are resilient to cyber attacks. However, consequences of CPS cybersecurity attacks can be less predictable and far more difficult to mitigate if cybersecurity threats were not adequately considered at the point of system design.
CPS owners and operators face a daunting challenge in keeping pace with the modern cybersecurity landscape. They find themselves at the intersection of digital technology and engineering and are often targeted in sophisticated cybersecurity attacks. Prominent cybersecurity attacks [3] on CPS have brought attention not only to the vulnerability of these systems, but also the inherent weaknesses of critical infrastructure reliant on them. These vulnerabilities could potentially result in severe consequences, affecting lives, economies, and national security.
To address these challenges, CPS owners and operators can take guidance from internationally recognised cybersecurity frameworks and standards. Firstly, the National Institute of Standards and Technology (NIST) Cybersecurity Framework [4] provides guidance to identify and manage cybersecurity risks in a systematic way. Secondly, the International Electrotechnical Commission (IEC) 62443-4-1:2018 standard [5] specifies process requirements for the secure development of products used in industrial automation and control systems, offering valuable guidance for securing CPS. These alone are not sufficient to mitigate ever-evolving cybersecurity challenges facing CPS. The dynamic threat landscape requires a tailored approach to CPS cybersecurity, combining continual assessment and adaptation of physical measures, cybersecurity practices, as well as risk mitigation controls.
To establish a clear framework for this paper, we define three key terms: “security modelling”, “threat modelling”,and “attack modelling”. Security modelling refers to the overall process of identifying, assessing, and mitigating risks to protect a system or organisation from a wide range of potential threats, both cyber and physical, throughout its entire lifecycle. Within this broader framework, threat modelling is a proactive approach used primarily in the early stages of a system’s lifecycle to anticipate potential cybersecurity threats and vulnerabilities, enabling the incorporation of relevant mitigation controls into the system design [6] [7] [8]. In contrast, attack modelling is a reactive approach that focuses on analysing specific attacker tactics, techniques, and procedures—including both cyber and physical actions—once the system is in operation, to develop targeted defenses and response strategies [9] [10] [11] [12]. To effectively mitigate risks in CPS, integrating threat and attack modelling into a continuous, iterative security modelling process is essential throughout the system’s lifecycle [13] [14].
In this paper, we design and conduct a systematic literature review (SLR) to study how existing security modelling, threat modelling and attack modelling approaches may be integrated in an iterative framework tailored to CPS in a continuous lifecycle context. This framework is crucial to ensure that CPS security measures are not static but evolve with the changing cybersecurity threat landscape. By bridging this gap, our survey aims to propose a holistic approach to enhance the cyber resilience of CPS.
Findings. In this review, we identified several state-of-the-art security modelling, threat modelling and attack modelling studies relevant to CPS, with the following findings:
  • Threat modelling is usually conducted in early stages of system development, and unlikely to anticipate new attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles.
  • Security models in the literature that focus on IT systems could be challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • There is limited differentiation between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks and consequences are considered in security modelling.
  • Unclear definitions and relationships between threat modelling and attack modelling can lead to inconsistent security modelling approaches. A unified security modelling approach that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article to address this.
The rest of the article is structured as follows: Section 2 provides an overview of the review’s background while Section 3 describes related work. Our methodology and results are outlined in Sections 4 and 5. Based on the results, we discuss our findings and research directions for future work in Section 6, and conclude in Section 7.
2 Background
This section provides some background into how CPS differ from conventional IT systems, and the characteristics of adversaries who target CPS in cyber attacks. Together, this explains why a tailored security modelling approach is necessary for CPS.
2.1 Cyber-Physical Systems
Different from traditional IT systems, a CPS is a System of Systems (SoS) that integrates computational elements, communication networks, and physical processes to form a unified system. CPS converges Information Technology (IT) and Operational Technology (OT) and has different priorities for cybersecurity. OT focuses on safety and reliability, while IT emphasises the confidentiality, integrity, and availability of information. Additionally, CPS networks must accommodate diverse communication modes, ranging from standalone to highly networked systems, potentially utilising legacy protocols such as serial communications, TCP/IP, or object exchange protocols. The heterogeneity and, in certain instances, expansive geographic scopes of CPS networks contribute to the complexity of modelling security for CPS.
Despite their longer system lifecycles, CPS are not upgraded or patched frequently as system or software changes may affect their reliability or even safety [15] [16] [17] [18]. This results in outdated software, exposing CPS to an ever-increasing number of cybersecurity vulnerabilities over time. At the same time, some CPS employ “security-by-obscurity”, using proprietary software and technology that may contain undisclosed security weaknesses that attackers could exploit. Moreover, the lack of “security-by-design” in CPS lowers the resilience to cybersecurity attacks even further, as CPS may lack security logs and fail to leverage log data for proactive cybersecurity monitoring and defence. Finally, there is a misconception that merely physically isolating CPS from external networks, known as an ”air gap” network security measure, is adequate to shield CPS from all cybersecurity threats. However, this notion is flawed, particularly in scenarios where multi-agent, multi-path CPS attacks could employ social engineering techniques, physical access, insider threats, and exploit supply-chain vulnerabilities, enabling attackers to breach the ”air gap” and compromise CPS, even when they are physically segregated from external networks.
2.2 Adversarial model
MITRE ATT&CK (Adversarial Tactics, Techniques, and Common Knowledge) [19] serves as a widely adopted knowledge base for characterising the actions and behaviours of cyber adversaries. Specifically tailored for ICS—a category of CPS—MITRE ATT&CK provides a matrix known as ”ATT&CK for ICS” [20]. This matrix focuses on the unique Tactics, Techniques, and Procedures (TTPs) employed by adversaries targeting industrial control environments. These TTPs play a crucial role in informing and guiding cybersecurity defence for CPS across the various stages of a cyber-attack lifecycle, as outlined in the Lockheed Martin cyber kill chain model [21].
An analysis of past cyber attacks on CPS reveals a taxonomy [22] highlighting differences from those observed in IT systems, necessitating a tailored security modelling approach for CPS. CPS attackers employ dynamic, multi-agent, and multi-path attack TTPs to defeat physical access controls and ”airgap” measures (isolating CPS from external networks). Additionally, adversaries evolve their TTPs over time, such as using ”low and slow” TTPs to evade detection and prolong access to compromised CPS.
2.3 CPS cyber-intrusion example
We present an example of a 6-month long CPS intrusion by a sophisticated attacker group comprising two adversaries A and B, adapted from one of the author’s professional experience. In this example, Adversary A compromises the organisation’s IT network to gain remote access to a CPS engineer’s computer and employs social engineering techniques to deliver custom malware to an isolated (”air gapped”) CPS network. Adversary B’s focus on the other hand is to develop custom malware to firstly profile the CPS, then to weaponise custom malware to manipulate connected Remote Terminal Units (RTUs) causing physical damage e.g. via actuators or valves. The diamond model of intrusion analysis [23] as well as a combination of the diamond model and the cyber kill-chain [24] are used to describe both the flow (Figure 1) and timeline (Figure 2) of this CPS cyber-attack.
Figure 1: CPS cyber intrusion flow representation using the diamond model of intrusion analysis
Figure 2: Timeline of a CPS cyber intrusion illustrating the steps in a cyber kill-chain
2.4 Security modelling framework for CPS
In the CPS cyber intrusion example, we demonstrated the key characteristics and inter-relationships between threat and attack modelling methods, namely the diamond model and cyber kill-chain respectively. Against this backdrop, we propose a unified security modelling framework for CPS (Figure 3) that integrates these modelling methods while drawing from continuous security monitoring across all stages of the CPS lifecycle for informed cybersecurity decision-making.
This security modelling framework for CPS does not replace threat modelling and attack modelling methods. Rather, the framework makes it easier and more practical for CPS cybersecurity practitioners to adopt and use these modelling methods to address evolving cybersecurity threats and attacker capabilities throughout the CPS lifecycle.
Figure 3: Security modelling framework for CPS that integrates threat modelling, attack modelling and security monitoring throughout the system lifecycle
3 Related Work
There are few existing systematic literature reviews on threat modelling or attack modelling. Humayed et al.’s literature survey [25] covered the security and privacy of cyber-physical systems, with a special focus on ICS, smart grids, medical devices, and smart cars. They presented a taxonomy of CPS threats, vulnerabilities, known attacks and existing controls, and importantly, captured how an attack of the physical domain of a CPS can result in unexpected consequences in the cyber domain and vice versa.
Xiong et al.’s SLR [8] was another early review on cybersecurity threat modelling, where one CPS threat modeling method was cited: Burmester et al.’s framework [26]. Based on the Byzantine paradigm, it enables the modeling of CPS security and facilitates formal analysis and security proof through cryptographic methodologies. Tatam et al. [27] reviewed threat modelling approaches for APT-style attacks, but the review did not focus on CPS security. The SLR by Khalid et al. [28] on the other hand, focused on threat modelling of industrial control systems (ICS). The authors described the various ICS threat models adopted in literature, and noted it was timely to consider a framework that covers all aspects of the various threat modelling methods.
Earlier, the evaluation of threat modelling methods by Shevchenko et al. [29] made a similar conclusion that there is no single method that can cover the full spectrum of CPS threats, and a framework that employed a combination of methods and techniques should be used. Supporting this argument, the hybrid threat model for smart systems proposed by Valenza et al. [30], and the work by Li et al. [31] on security-usability threat modelling for industrial control systems underscored the need to consider not only cyber but also human factors in ICS threat modelling due to the system-of-systems (SoS) nature of CPS.
These CPS threat modelling methods and techniques were well-covered in the literature, with case studies in specific industries and domains, e.g. the study of smart grid CPS cybersecurity by Nafees et al. [32], the study of cyber-physical energy systems security (CPES) by Zografopoulos et al. [33], the application of the STRIDE threat model in modern vehicle vulnerability assessments by Abuabed et al. [34], the work on CPS threat modelling for power transformers by Ahn et al. [35], the work on medical CPS threat modelling by Almohri et al. [36], the work by Al Asif et al. [37] on telesurgery system threat modelling, the work by Jbair et al. [38] on threat modelling for industrial cyber-physical systems (ICPS) in the manufacturing industry, and the work on threat modelling in smart firefighting systems by Zahid et al. [39].
4 Methodology
We followed the guidelines for systematic literature reviews proposed by Kitchenham et al. [40]. The process of our literature review is shown in Figure 4.
Figure 4: Systematic literature review process
4.1 Research questions
This review aims to answer the following Research Questions (RQs):
  • RQ1: What CPS threat or attack models have been adopted in existing literature?
    The rationale for this research question is to understand what the “state-of-the-art” research in this area is.
  • RQ1.1: How do the proposed threat and attack models align with cybersecurity in various phases of the CPS lifecycle?
    The rationale for this research question is to understand how existing CPS threat and attack models address cybersecurity threats not only in the design phase, but when CPS are in operational phases. This is particularly crucial as CPS may operate for many years, and likely render threat models considered in the system design irrelevant.
  • RQ1.2: In what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?
    The rationale for this research question is to understand how existing CPS threat and attack models are augmented and improved with real-world feedback from cybersecurity intrusions, near-misses, or external threat intelligence information. This is an important bridge between theoretical models and practice, that can lead to tangible security improvements in CPS.
  • RQ2: What research gaps have been identified in the methods used for modelling threats and attacks in CPS?
    The rationale for this research question is to help identify research gaps in current CPS security modelling methodologies and to direct future research efforts.
4.2 Search protocol
This sub-section outlines the search protocol used in this review.
We used both US and UK English variations for all search terms. For example, we included both ”modelling” (UK English) and ”modeling” (US English).
The databases selected for our search were the IEEE Computer Society Digital Library, Science Direct, ACM Digital Library, and Scopus. These databases were chosen because they are recognised as reputable and reliable repositories for academic literature and have an extensive collection of articles and publications related to cybersecurity.
While multidisciplinary approaches that integrate principles from control systems engineering, safety engineering, and human factors are vital for addressing the intricate interplay between cyber and physical components in CPS environments, they fall outside the scope of this paper. A comprehensive review of the full range of strategies required for effective CPS protection would necessitate insights from various specialised domains, each deserving dedicated attention and expertise.
Our primary focus was therefore on search terms related to cybersecurity modelling, security modelling, threat modelling, and attack modelling, in alignment with the subject of our literature review. We defined Criteria A as our initial list of search keywords, which included: ”security model,” ”security modelling,” ”cybersecurity model,” ”cybersecurity modelling,” ”threat model,” ”threat modelling,” ”attack model,” and ”attack modelling.” We incorporated both ”model” and ”modelling” to maximise the coverage of relevant search results.
As we wanted to narrow our search results to articles on security, threat, or attack modelling, we limited the initial search to article titles. Minor adjustments to the search terms were necessary, depending on the Boolean vocabulary supported by each database.
  • At the time of writing, the IEEE Computer Society Digital Library does not support in-line Boolean operators like ”OR” and ”AND” in the search term input. Therefore, we used the Advanced Search menu, selected “in Document Title,” and conducted repeated searches using individual search keywords, such as ”security modelling” and ”attack model”.
  • On ScienceDirect, we selected Advanced Search and entered (“cybersecurity model” OR “security model” OR “threat model” OR “attack model”) in the Title field. The Boolean “OR” had to be specified, and there was no need to specify the “modeling” keyword as it is implied by “model” in the search term.
  • On ACM Digital Library, we selected “Advanced Search”, selected “Title” under “Search Within”, then entered: ”security model” OR ”security modeling” OR ”security modelling” OR ”cybersecurity model” OR ”cybersecurity modeling” OR ”cybersecurity modelling” OR ”threat model” OR ”threat modeling” OR ”threat modelling” OR ”attack model” OR ”attack modeling” OR ”attack modelling”.
  • On Scopus, we selected “Article title” under the “Search within” field, then entered ”security model” OR ”security modeling” OR ”security modelling” OR ”cybersecurity model” OR ”cybersecurity modeling” OR ”cybersecurity modelling” OR ”threat model” OR ”threat modeling” OR ”threat modelling” OR ”attack model” OR ”attack modeling” OR ”attack modelling” in the “Search documents” field.
We then refined the search results further by specifying additional keywords: (”cyber physical,” CPS, SCADA, ICS, APT), to search within article contents. We included SCADA (Supervisory Control and Data Acquisition), ICS (Industrial Control System), and APT (Advanced Persistent Threat) to ensure that the search results were relevant to the adversarial tactics, techniques, and procedures associated with CPS cyber attacks. This list of search terms constitutes Criteria B.
  • On the IEEE Computer Society Digital Library, we used the Advanced Search menu with the initial keywords, selected the “AND” and “in Document Content” options, and entered the additional keywords. We noted that there might be duplicate search results due to the use of different additional keywords, which needed to be identified and removed.
  • On ScienceDirect, we selected Advanced Search and entered (”cyber physical” OR CPS OR SCADA OR ICS OR APT) in the “Find articles with these terms” field.
  • On ACM Digital Library, we selected “Edit Search”, added a new search field, selected “Full text” under “Search Within”, then entered: ”cyber physical” OR CPS OR SCADA OR ICS OR APT.
  • On Scopus, we selected “Add search field”, selected “Article title, Abstract, Keywords” under the “Search within” field, and entered: ”cyber physical” OR CPS OR SCADA OR ICS OR APT.
In the final phase of the search protocol, we removed repeated articles from the combined search results based on similar article titles. We also excluded articles that were not related to the research questions, after reviewing their titles and abstracts.
4.3 Categorisation and classification
We categorised the selected articles into three distinct clusters, based on their primary focus and content:
  • Cluster 1 (C1): Articles focused on threat modelling method(s)
    Articles are placed in this cluster if their primary focus is on threat modelling, defined as the process of identifying potential threats and vulnerabilities early in the system lifecycle to inform the design of mitigation strategies. These articles predominantly discuss frameworks, tools, techniques, or methodologies specifically designed for threat modelling.
  • Cluster 2 (C2): Articles focused on attack modelling method(s)
    Articles are included in this cluster if they primarily focus on attack modelling, which involves analysing attacker tactics, techniques, and procedures (TTPs) to inform defence strategies during the system’s operational phase. These articles are centered on understanding the behavior, capabilities, and goals of attackers and often provide detailed descriptions of attacker models, simulation methods, or real-world case studies of cyber attacks.
  • Cluster 3 (C3): Articles that are systematic literature reviews (SLRs)
    Articles are categorised in this cluster if they are systematic literature reviews that provide a comprehensive review of existing threat and/or attack modelling methods. These articles systematically analyse and synthesise findings from a range of studies, offering insights into the state of research, common practices, gaps, and future directions.
Articles were categorised based on a thorough analysis of their content, including their primary objectives, methodologies, and the specific aspects of threat or attack modelling they emphasise. Where articles covered both threat and attack modelling, they were classified according to the dominant theme or focus, based on the extent and depth of discussion related to either method.
Additionally, articles were categorised according to the following criteria:
  1. 1.
    Year of publication
  2. 2.
    Document type, i.e. conference paper, article, review, book chapter, etc.
  3. 3.
    Number of citations (based on Google Scholar Citations)
  4. 4.
    Research question(s) article is relevant to
4.4 Quality assessment
In this phase, we reviewed the articles to ensure they were relevant to CPS security, threat or attack modelling, and we removed articles that do not describe these models’ approaches, CPS applications, or CPS case-studies. We also reviewed bibliographies in the articles and included additional articles that may be relevant to the review.
4.5 Data collection
We studied the articles closely to understand the background, approaches, and results from the various research. Importantly, we sought to identify gaps in the research in terms of security modelling in the context of CPS cyber attacks. We documented our notes from the review in a public repository [41] for future reference.
4.6 Data analysis
In the final phase of the SLR methodology, we synthesised our findings and documented our analysis as well future research directions.
5 Results
The search was conducted on the following digital libraries in August 2024: IEEE Computer Society Digital Library, Science Direct, ACM Digital Library and Scopus. Following the search, we had 366 articles in total. We performed a rigorous quality assessment, where we first excluded duplicate articles, then we identified and excluded irrelevant articles based on their abstracts. We then performed full-reading of about 32 articles that remained and excluded articles that were not relevant to any of our research questions. Finally, we had 15 articles subject to further analysis (Figure 5).
Figure 5: Literature survey search results (August 2024)
The articles were categorised into three clusters (Table 1) based on full reading of the articles: threat modelling (C1), attack modelling (C2), and SLRs (C3). There were no articles on “security modelling” that were relevant to CPS.
Table 1: Categorisation of CPS cybersecurity articles
Reference Title Threat Modelling (C1) Attack Modelling (C2) SLRs (C3)
Ayrour et al. (2018) Modelling cyber attacks: a survey study $\bullet$
Fernandez (2016) Threat modeling in cyber-physical systems $\bullet$
Khalil et al. (2022) Threat Modeling of Cyber-Physical Systems-A Case Study of a Microgrid System $\bullet$
Khalil et al. (2023) Threat modeling of industrial control systems: A systematic literature review $\bullet$
Khan et al. (2017) STRIDE-based threat modeling for cyber-physical systems $\bullet$
Kim et al. (2022) STRIDE‐based threat modeling and DREAD evaluation for the distributed control system in the oil refinery $\bullet$
Kumar, R., et al. (2022) APT attacks on industrial control systems: A tale of three incidents $\bullet$
Martins et al. (2015) Towards a systematic threat modeling approach for cyber-physical systems $\bullet$
Mekdad et al. (2021) A threat model method for ICS malware: the TRISIS case $\bullet$
Neubert and Vielhauer (2020) Kill chain attack modelling for hidden channel attack scenarios in industrial control systems $\bullet$
Paudel et al. (2017) Attack models for advanced persistent threats in smart grid wide area monitoring $\bullet$
Tatam et al. (2021) A review of threat modelling approaches for APT-style attacks $\bullet$
Xiong and Lagerström (2019) Threat modeling – A systematic literature review $\bullet$
Yang and Zhang (2023) From Tactics to Techniques: A Systematic Attack Modeling for Advanced Persistent Threats in Industrial Control Systems $\bullet$
Zalewski et al. (2013) Threat modeling for security assessment in cyberphysical systems $\bullet$
There were 10 articles in C1 and five in C2, showing that existing literature delved into threat modelling methods more than attack modelling methods. Next, we then classified the articles based on the criteria described in Section 4.3.
The first selected article was published in 2013 and although no article was published in 2014, the number of relevant articles has gradually risen over the years (Figure 6), signaling a growing level of research interest in CPS security modelling. Furthermore, out of the 15 selected articles, seven were from conference proceedings and eight were from journals. This indicated that almost an equal number of articles in CPS cybersecurity modelling were published in journals and conferences.
Figure 6: Number of CPS cybersecurity articles per year
Next, we used citation metrics on Google Scholar to identify articles that had more than 100 citations. As of January 2024, only two of the 15 selected articles had more than 100 citations (Table 2). The top cited article was on STRIDE-based threat modelling for cyber-physical systems by [42], followed by the SLR on threat modelling by [8]. This observation, together with that from the number of articles published yearly, suggested that this field of research is still relatively new.
Table 2: Top-cited CPS cybersecurity articles
Reference Citations
Khan et al. (2017) 239
Xiong and Lagerström (2019) 231
Finally, we performed full reading of the 15 articles and analysed which research questions they were relevant to (Table 3).
Table 3: Article relevance to research questions
Reference Cluster RQ1 RQ1.1 RQ1.2 RQ2
Fernandez (2016) C1 $\bullet$
Khalil et al. (2022) C1 $\bullet$ $\bullet$
Khan et al. (2017) C1 $\bullet$
Kim et al. (2022) C1 $\bullet$ $\bullet$
Martins et al. (2015) C1 $\bullet$
Mekdad et al. (2021) C1 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Zalewski et al. (2013) C1 $\bullet$ $\bullet$
Kumar et al. (2022) C2 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Neubert & Vielhauer (2020) C2 $\bullet$ $\bullet$ $\bullet$
Paudel et al. (2017) C2 $\bullet$ $\bullet$ $\bullet$
Yang and Zhang (2023) C2 $\bullet$ $\bullet$ $\bullet$
Ayrour et al. (2018) C3 $\bullet$ $\bullet$ $\bullet$ $\bullet$
Khalil et al. (2023) C3 $\bullet$ $\bullet$
Tatam et al. (2021) C3 $\bullet$ $\bullet$
Xiong and Lagerström (2019) C3 $\bullet$
Three of the articles [43] [3] [44] addressed all the research questions. The remaining articles addressed some, but not all the research questions. 12 of the 15 articles were relevant to RQ1, and only two of the seven C1 articles were relevant to RQ1.1 (28.57%), one was relevant to RQ1.2 (14.29%), and three (42.86%) were relevant to RQ2. On the other hand, three of the four C2 articles were relevant to RQ1.1 (75%), four were relevant to RQ1.2 (100%) and two were relevant to RQ2 (50%). Lastly, three of the four C3 articles were relevant to RQ1.1 (75%), one was relevant to RQ1.2 (25%), and four were relevant to RQ2.
A comparison of articles across the clusters (Figure 7) showed that existing literature that focused on threat modelling (C1) addressed RQ1 more than the other research questions. Those that were focused on attack modelling (C2) appeared to address RQ2 less. Lastly, the C3 articles did not address RQ1 as they were literature reviews, and these articles appeared to address RQ1.2 less.
Figure 7: Relevant articles across clusters
6 Discussion
This section aims to describe the findings and limitations of this review, as well as future research directions.
6.1 Findings
The selected 15 articles were categorised into three clusters: threat modelling methods (C1), attack modelling methods (C2), and SLRs (C3). There were eventually seven C1 articles, five C2 articles, and three C3 articles.
6.1.1 RQ1
To address the first research question, “What CPS threat or attack models have been adopted in existing literature?”, we examined the threat and attack modelling methods described in the articles, excluding the SLRs. SLRs were excluded because they primarily focus on summarising and synthesising existing research rather than presenting original threat or attack modelling methods. Therefore, they provide an overview of methodologies rather than specific models. The methods we identified include: the Diamond Intrusion Analysis model [23], attack trees [10] [11], the Lockheed Martin cyber kill chain [21], and the MITRE ATT&CK framework [19], as detailed in Table 4.
Table 4: Threat and attack modelling methods adopted in selected articles
Reference Cluster Threat and Attack Modelling Method(s)
Fernandez (2016) C1 Misuse patterns
Khalil et al. (2022) C1 STRIDE
Khan et al. (2017) C1 STRIDE
Kim et al. (2022) C1 STRIDE, DREAD
Martins et al. (2015) C1 Generic Modeling Environment (GME)
Mekdad et al. (2021) C1 Diamond Model of Intrusion Analysis
Zalewski et al. (2013) C1 STRIDE
Kumar et al. (2022) C2 Attack tree
Neubert and Vielhauer (2020) C2 Kill chain
Paudel et al. (2017) C2 Attack tree
Yang and Zhang (2023) C2 Abstract APT attack model
6.1.2 RQ1.1
On the second research question, “how do the proposed threat and attack models align with cybersecurity in various phases of the CPS lifecycle?”, we found that threat modelling was conducted in the early stages of the system development cycle [45], or more specifically, during system design and validation phases [42] [46] [47]. Interestingly, unlike threat modelling, none of the C2 articles suggested or mentioned which phase of the system lifecycle attack modelling should be performed in. A possible reason is that attack modelling is seen as a complement to threat modelling methods [45] [3].
6.1.3 RQ1.2
On the third research question, “in what manner do the proposed threat and attack models adjust to evolving attacker tactics, techniques, and procedures over time?”, we found differing views on how to incorporate attacker tactics, techniques, and procedures (TTPs) into threat modelling frameworks. For example, Khalid et al. [45] argued that attack categories in their proposed taxonomy should be abstract and should not include specific details of attack techniques, such as those outlined in MITRE ATT&CK for Industrial Control Systems [20]. In contrast, Fernandez et al. [48] proposed a threat model using misuse patterns that may include attacker TTPs to represent CPS threats. These differing approaches highlight the challenge of balancing model specificity with adaptability to evolving threats.
Several studies have attempted to incorporate attacker TTPs or similar behavioural characteristics into their models. For instance, Mekdad et al. [44] combined the Diamond Model of Intrusion Analysis [23] and the ICS kill chain [49] to develop a dynamic ICS threat model. Similarly, attack trees [11] have been employed to model CPS attacks [3] [14], while Neuber et al. [50] demonstrated how the Lockheed Martin Cyber Kill Chain [21] could be used to model hidden channel attack scenarios in ICS. Yang et al. [13] proposed an abstract APT attack model capable of recognising and correlating attacker TTPs to identify and mitigate complex attacks in ICS. This model emphasises the need to consider a broad range of TTPs to reflect the evolving threat landscape accurately.
However, our analysis revealed that while some models incorporate elements of adaptability to evolving TTPs, there is a significant gap in addressing how these models dynamically adjust to the continuously changing tactics of sophisticated adversaries. For example, while models like those proposed by Mekdad et al. and Yang et al. suggest mechanisms for integrating new TTPs, they do not specify processes for continuously updating or validating these models against emerging threats.
Moreover, while several articles acknowledged the need for models to account for changes over time, they did not explicitly address attacker TTPs as a component of these changes. For instance, Almohri et al. [36] noted that the threat model of a medical CPS might change with its corresponding trust model, incorporating time-bound elements such as temporarily trusted individuals or accounts. Similarly, Khalil et al. [45] highlighted that systems may evolve over time, necessitating the inclusion of previously excluded information in future threat models to account for system modifications.
Our findings raise the question of whether threat modelling continues to be relevant in CPS security design if they are conducted in the early stages of system development, given the usually long periods of development and implementation for complex CPS infrastructure. The cyber threats and available technologies to defend against those threats would have changed by the time the CPS become operational, often resulting in insecure systems. Relatedly, this presents a critical issue potentially affecting national security. CPS are usually used in critical national infrastructure and for that reason, are a prime target of sophisticated adversaries who have extended resources, capabilities and experience, may adopt “low and slow” tactics to infiltrate and maintain persistence in compromised CPS over many months and years. Outdated threat models would not be able to adjust and enable swift detection and eradication of these persistent threats.
This gap suggests a need for more dynamic and resilient security models that can adapt to both the evolving threat landscape and the changing nature of CPS environments.
6.1.4 RQ2
For our final research question, “what research gaps have been identified in the methods used for modelling threats and attacks in CPS?”, we made several observations from the literature.
Security models mostly focus on IT systems. Firstly, threat and attack models in the literature mostly focus on IT systems, even in the context of CPS security. For example, the STRIDE model developed by Microsoft was used in several articles to identify CPS threats, notwithstanding that the model is more commonly used in the context of software development and system security. Using STRIDE in CPS threat modelling [34] [42] [46] has several limitations which were not obvious in the literature.
Specifically, the STRIDE approach relies on Data Flow Diagrams (DFDs) to identify system entities and trust boundaries, which has several shortcomings related to inadequate representation of security concepts, data elements, abstraction levels, and deployment information [51]. Moreover, it is sometimes not practical or even possible to construct an accurate DFD in CPS. There are various reasons for this. A CPS is a system-of-systems that includes both cyber and physical components. Availability, accuracy and integrity of data flows between cyber and physical components cannot be assumed. Most importantly, the consequences of CPS compromise extend beyond those specified in the STRIDE model, e.g. physical changes or effects.
This limitation may be addressed by leveraging diverse, multi-disciplinary expertise from physical security and domain specialists to augment the DFD by capturing interdependencies and interactions between cyber and physical components, and not be limited to IT systems. This limitation may also be addressed by continuously refining the DFD based on real-time data and feedback from system operations and cybersecurity events detected in the CPS environment over time.
Relatedly, the literature refers to CPS use-cases in the context of ICS, smart grids, medical CPS (MCPS), SCADA systems, or modern vehicles. We expect that CPS use-cases will increasingly include new scenarios involving technologies such as robotics, drones, EV charging infrastructure and smart buildings. Conventional threat modelling approaches, primarily tailored for IT systems, may encounter heightened challenges when addressing cybersecurity threats in these emerging domains.
Attack scenarios are not realistic. Secondly, most threat and attack modelling methods in the literature made assumptions based on simplistic cybersecurity attack scenarios, and did not consider the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS. The exceptions were: Fernandez et al. [48] who proposed a patterns-based threat model that considers multiple vulnerabilities exploited in different parts of a distributed system, and Zografopoulos et al. [33] who analysed Cyber-Physical Energy Systems (CPES) security from a multi-layered attack perspective.
Differentiation between IT and CPS cybersecurity breaches. Thirdly, most of the literature did not differentiate cybersecurity breaches in CPS from those in IT systems in that they are not a binary (all or nothing) event. Instead of abrupt, catastrophic system failures, CPS cybersecurity breaches often result in gradual degradation of system services before the CPS fails eventually [52]. This is analogous to a system fault in an autonomous vehicle, the vehicle does not come to a stop immediately—which would injure the passengers—but instead gradually slows to a safe stop. The implication to CPS security is that from a cyber defence perspective, there is an opportunity to remediate intrusions through cyber or physical interventions [24] or to perform dynamic reconfiguration and self-healing tasks [33] before the CPS fails completely.
Consequence-driven considerations. CPS increasingly face cybersecurity threats that can endanger the safety, reliability and resilience of critical infrastructure. With the convergence of cybersecurity, safety, and reliability in CPS, it would be beneficial to adopt a consequence-driven and cyber-informed approach [53] towards threat and attack modelling, and align with system development processes in engineering domains such as the V-model that is commonly used in the automotive industry [54], as well as the failure mode and effect analysis (FMEA) and fault-tree analysis (FTA) techniques used in safety and reliability engineering domains [16]. Doing so will avoid making threat or attack modelling a one-time event that assumes static cybersecurity threats and vulnerabilities, resulting in inherent security weaknesses that are “baked” into proprietary CPS systems and software designs.
Distinguishing between CPS threat and attack modelling. Lastly, we found that while the literature used “threat modelling” and “attack modelling” interchangeably, both terms usually referred to threat modelling and not the latter. On the other hand, some authors promoted definitions of threat modelling from literature and distinguished threat from attack modelling [28] [8], while others proposed security modelling frameworks that are based on threat models [26]. In this paper, we provide explicit definitions for both threat and attack modelling to clarify their distinct roles within cybersecurity (see Section 1). Based on these definitions and the literature, we analysed the various characteristics and inter-relationships between threat and attack modelling methods and these are aligned with our proposed unified security modelling framework for CPS as described in Section 2.4.
6.2 Future work
The limitations identified in the literature can be summarised as follows: existing security models predominantly focus on IT systems, neglecting the unique complexities of CPS scenarios and overlooking the multi-layered, multi-path, and multi-agent nature of real-world CPS attacks. Additionally, they often do not distinguish the consequences of cybersecurity breaches in CPS from those in IT systems and tend to treat cybersecurity threats as static and one-time events, contributing to inherent security weaknesses embedded within proprietary CPS systems and software designs.
Against this backdrop, our future work may be to research cybersecurity anomaly detection in CPS, to review the extent to which real-world cybersecurity intrusions and threat intelligence can inform and at the same time, enhance security modelling throughout the CPS system lifecycle. A more context-driven security model as this would better enable the identification of vulnerabilities, understanding of attack propagation mechanisms, and evaluation of system resilience.
We may also use Petri net based approaches [55] [56] to model CPS intrusions better by integrating both cyber and physical actions. Petri nets are a graphical and mathematical tool for studying concurrent and distributed systems and have been widely applied in many different areas of computer science and other disciplines [57]. In the context of CPS intrusion analysis, Petri nets are invaluable for modelling the interactions between cyber and physical components over time, thereby overcoming limitations of existing security modelling approaches when considering dynamic, multi-agent cyber-physical threats.
Finally, our future work may be to research self-healing techniques in CPS, to bolster existing cybersecurity defences and strengthen system resilience whether through cyber or physical actions, against sophisticated adversaries and threats.
6.3 Limitations of this review
Although we aimed to cover as much existing literature on security, threat and attack modelling in CPS as possible, we limited our search protocol to leading scientific databases as the quality of the articles may be ensured. Furthermore, we limited our search to journals and conference proceedings and consequently, we may miss out relevant articles from other sources such as book chapters, etc.
We also limited the initial scope of search to keywords in the article titles. This approach may have excluded articles relevant to security modelling, threat modelling, or attack modelling that did not explicitly mention these terms in their titles. However, to partially mitigate this exclusion, we reviewed the bibliographies of the selected papers and added any other relevant articles that may have been missed initially. Nevertheless, we ensured to select only articles that are pertinent to the research questions for eventual review. Based on the review of abstracts of the shortlisted articles from the search, we believe the coverage of the review to be sufficiently comprehensive.
Our research questions were focused on how existing security modelling approaches addressed CPS, and corresponding adversarial tactics, techniques, and procedures. These factors may introduce some bias to the articles selected for review and may miss out articles that are relevant to CPS but did not mention adversarial models specifically. We addressed this by reviewing bibliographies in the selected articles and then including additional articles that may be relevant to the review.
7 Conclusion
In this review, we identified several state-of-the-art threat modelling and attack modelling studies relevant to CPS, with the following findings:
  • Threat modelling is usually conducted in early stages of system development, and unlikely to anticipate new attacker tactics, techniques and procedures (TTPs) in later stages of system lifecycles.
  • Security models in the literature that focus on IT systems could be challenging to use when modelling CPS cybersecurity threats and attacks. This is a pertinent issue for practitioners given the multi-layer, multi-path or multi-agent characteristics of real-world cybersecurity attacks in CPS.
  • There is limited differentiation between cybersecurity breaches in IT systems versus those in CPS. Unlike IT systems, cybersecurity incidents in CPS can result in complex failure modes, as well as consequences in both cyber and physical domains. Adopting a consequence-driven and cyber-informed approach to CPS security is vital towards ensuring that cyber and physical attacks and consequences are considered in security modelling.
  • Unclear definitions and relationships between threat modelling and attack modelling can lead to inconsistent security modelling approaches. A unified security modelling approach that integrates threat modelling, attack modelling, and security monitoring to enhance the cyber resilience of CPS is proposed in this article to address this.
While the number of published articles is comparatively small compared to more established research domains, there is visible growing interest in CPS cybersecurity research, as evidenced by the rising number of publications since 2013. Anticipating the broader use of CPS beyond traditional industrial scenarios, we envisage continued improvements in both the diversity and quality of articles in this domain, underscoring the growing significance of CPS cybersecurity.
Declarations
The authors make the following declarations about this work:
Availability of data and materials. This article does not involve the generation of any original datasets or materials. All data and resources referenced in this study are publicly available and can be accessed through the sources cited within the article.
Competing interests. The authors have no competing interests to declare.
Funding. The authors did not receive support from any organisation for the submitted work.
Authors’ contributions. All authors contributed to the study conception and design. The literature review and analysis were performed by Shaofei Huang. The first draft of the manuscript was prepared by Shaofei Huang and all authors reviewed and edited previous versions of the manuscript. All authors read and approved the final manuscript.
Acknowledgements. The authors have no acknowledgments to declare.